Session: Oct 2008 (session 3)
Wavelength: 0.7 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Michael Lindqvist at Thu Oct 16 16:09:46 2008) Success :
Some rain last hour(s). Otherwise no known problems.

Effelsberg (Uwe Bach at Mon Oct 20 07:30:04 2008) Minor failures :
Success, but a few scan were lost for pointing. It was raining at the beginning but cleared up later.

Noto (P. Cassaro at Fri Oct 31 09:57:31 2008) Success :